[Search for users] [Overall Top Noters] [List of all Conferences] [Download this site]

Conference asdg::cmp

Title:CMP
Notice:Welcome to the CMP notesfile
Moderator:FABSIX::B_TOWERLLE
Created:Tue Oct 31 1995
Last Modified:Fri Jun 06 1997
Last Successful Update:Fri Jun 06 1997
Number of topics:94
Total number of notes:6381

15.0. "STRASBAUGH GENERAL ISSUES" by SUBPAC::BJUBINVILLE (CMP-6 Equipment Engineering) Sat Sep 28 1996 04:30

    
     
        This note is designated for the calibration variables,
      recipes and other data files on the Strasbaugh tools for
      reference or accidental deletion.
    
                                                       BJ
    
T.RTitleUserPersonal
Name
DateLines
15.1slurry lossLUDWIG::RIDLONEliminate the obliviousSun Sep 29 1996 04:5911
    Both tools crashed tonight (A1 @1, A2 @1:30) as a result of loosing slurry
    to the room. Ahsland was called and responded in a timely manner. His
    report was that the filter had clogged (which was changed on Friday).
    Two product wafers were lost per tool.
    
    			
    						B-shift
    
    
    
    						
15.2particlesSTRATA::RIDLONEliminate the obliviousTue Oct 22 1996 00:1611
    All tools failed for particles. Tim Dyer called in and requested that
    the slurry day tank be flushed and cleaned, a new filter installed,
    and, if possible, change slurry lot number. Ahsland was notified and
    could do the first two but we don't have any other lot numbers in
    house. Once they were done we flushed slurry through the lines using
    for about 5 minutes and then the lines were di flushed for 10. The pad
    was then scrubbed and mqc's requested.
    
    
    					B-shift
    
15.3filter changeSTRATA::FINNFri Oct 25 1996 17:3729
The Fab6 CMP slurry system has been reconnected to the Osmmonic Filter 
system. Ashland Chemical has agreed to replace the filters twice this weekend 
to preventive filter clogging and the subsequent loss of slurry. Due to the 
current filter configuration, it will be necessary to shut the CMP room 
DOWN for a period of approximately one hour during each of these two filter 
change events. 

	The First filter change will occur tomorrow at 4:00PM. 

	The Second filter change will occur Monday morning at 6:30 AM.

Please ensure that all processing has stopped prior to these times. In 
addition Ashland Technicians will call the room at ext:4774 to confirm that 
processing on the tools has ceased prior to their beginning work.

Ashland Technicians will again call the room when their work has been
completed and we may continue processing. Their primary contact will be the 
operation group at ext 4774. In the event they are unable to reach anyone in 
the room, the following names have been given to Ashland as backup contacts:

	 John Swenson (pager 697) for Saturday afternoon 4:00pm
	 Mitch Ridlon (pager 652) for Monday morning 6:30 am

Thank you for your anticipated cooperation. 


Brendan 

15.4fitler updateLUDWIG::RIDLONEliminate the obliviousSun Oct 27 1996 21:052
    Slurry room checked at 2015. Input pressure to filter at 26 output at
    22. Flow good, pumping good. Will keep an eye on it.
15.5filter checkLUDWIG::RIDLONEliminate the obliviousMon Oct 28 1996 04:122
    Checked slurry room at 0245. Pressure looking good at 26 in 22 out. 
    Flow looking good and pumping souds steady. 
15.6filter checkLUDWIG::RIDLONEliminate the obliviousMon Oct 28 1996 05:512
    Checked slurry room at 0430. Pressure maintaining at 26 in 22 out.
    Flow's good, pump still pumping.
15.7filter checkLUDWIG::RIDLONEliminate the obliviousMon Oct 28 1996 06:343
    Checked slurry room at 0630. Pressure starting to fall off. 24 in 19
    out. Flow's still good. Pump starting to work harder. Talked with
    Ashland dudes; eight filters due in some time today.
15.8power glitchFABSIX::M_RIDLONEliminate the obliviousMon Nov 25 1996 07:077
    Power glitch @0645. B tools not running. A tools were. Didn't lose any
    wafers but had to power down tools to reset the motor controllers. A1
    had product but they should be able to touch them up. A2 just had
    50k's. 
    
    
    						B-shift
15.9Gimbal Bearings Replaced - #3 & #11ASDG::POIRIERWed Jan 22 1997 19:236
    
    Carrier Gimbal Assemblies Rebuilt:
    ---------------------------------
    
    Gimbal Assemblies #3 and #11 have had their gimbal bearings replaced.   
                      
15.10gimbal assy'sFABSIX::J_SWENSONThu Jan 23 1997 19:339
     Carrier Gimbal Assy's Rebuilt
     ----------------------------
    
    Carriers 1 and 13 had there gambal assys rebuilt
    Had problem with carrier # 2 will look at tomorrow.
    
    
    
                                       C-Shift 
15.11Bad Pads Removed From RoomASDG::POIRIERSat Apr 12 1997 20:469
    
    Pad Problems:
    ------------
    Rodel IC 1000 pad lot # 6HO-JP2 has been removed from the room due
    to repeated poor performance. We will leave them in Jim Tooker's
    office.
    
    
                                                              J.P.